OpenStudy (anonymous):

can someone check my work? tell me what answer you got for this problem (I got 1): Evaluate when a=-4 b=2 and c=-3 (-c+(1/2)(bc/\2-(bc)/\2)+ radical (abc+1)) + 1

OpenStudy (austinl):

\(-c+(\frac{1}{2})(bc^2-(bc)^2)+\sqrt{abc+1})+1\) Is this correct?

OpenStudy (anonymous):

precisely!

OpenStudy (austinl):

Ok, you just plug in the values and calculate! a=-4 b=2 and c=-3 \((-(-3)+(\frac{1}{2})((2)(-3)^2-((2)(-3))^2)+\sqrt{(-4)(2)(-3)+1}~)+1\)